how do you slove r+1/3=8/3
Found 3 solutions by checkley77, rfer, Alan3354:
Answer by checkley77(12844)

r+1/3=8/3 subtract 1/3 from both sides.
r=8/3-1/3 combine the fractions.
r=7/3 ans.

Answer by rfer(16322)

r+1/3=8/3
r=8/3-1/3
r=7/3 or 2 1/3

Answer by Alan3354(69443)